OpenStudy (anonymous):

A group of 11 people are going to the movies one Saturday. Adult tickets cost $11.50 each and child tickets cost $7.25 each. If the total cost for admission to the movies is $105.25, how many children attended the movie? a. 4 b. 5 c. 6 d. 7

OpenStudy (anonymous):

let the no of children be x then the number of adult in grp of 11 is 11-x as per the question we have 7.25x+11.50(11-x) =105.25 now solve for x ...

OpenStudy (anonymous):

you cant solve this

OpenStudy (anonymous):

u should try to solve it

OpenStudy (anonymous):

i can help u in solving..

OpenStudy (anonymous):

can you walk me through the steps? i just tried solving

OpenStudy (anonymous):

first simplify the left side

OpenStudy (anonymous):

18.75(11-x)

OpenStudy (anonymous):

so we have 7.25x+11.50(11-x) =105.25 7.25x +126.5 -11.5x =105.25

OpenStudy (anonymous):

7.25x -11.5x =105.25 -126.5 -4.25 x = -21.25 4.25x =21.25 x= 21.25/4.25 x=5

OpenStudy (anonymous):

ugh thank you i still dont get some of it but illl study a little more!

OpenStudy (anonymous):

yw which portion u feel confused..

OpenStudy (ybarrap):

Two equations. Two unknowns. Nice combination: Let A = Number of Adults attending Let C = Number of Children attending Total = A+C=11 people attending (this is one equation) \(A\times 11.50 + C \times 7.75 = 105.25\) (our second equation) Now, using the first we can solve for either C or A, let's go with C: C = 11-A. Now, stick this into the second equation. We'll then have one equation. One unknown. The best. \(A\times 11.50 + (11-A) \times 7.75 = 105.25\) Let's simplify. We need to bring the terms that go with \(A\) together and swing those constants to the other side of our equation: \(\large A(11.50-7.75)=105.25-11\times7.75 \implies A={105.25-11\times7.75\over 11.50-7.75}\). We now have A. Stick this into our first equation -- another equation with with just one unknown (after subbing). And we're done!